### v1

@@ Abstract
Spectral Conjugate Gradient Method Dai-Liao Method Optimization Signal Processing Secant Condition
MSC65K05 90C30 49M37
@@ 1 Introduction
where the stepsize $\alpha_{n}$ is a positive real parameter determined via exact or inexact line search, $x_{n}$ is the previous iterative point, $x_{n+1}$ is the current point, and $d_{n}$ is a suitable search direction. The class of algorithms that follow the form (2) is known as line search algorithms. For such algorithms, only the search direction $d_{n}\in\mathbb{R}^{m}$ and the stepsize $\alpha_{n}\in(0,\infty)$ are required. An appropriate descent direction $d_{n}$ must be chosen such that it satisfies the descent condition:
@@
Inspired by Shanno’s matrix perspective on CG algorithms [10], Babaie-Kafaki and Ghanbari [11] observed that the Dai-Liao search directions can be expressed as $d_{n+1}=-P_{n+1}\nabla f(x_{n+1})$ for all $n\in\mathbb{N}_{0}$, where $P_{n+1}=I-\frac{s_{n}y_{n}^{T}}{\langle s_{n},y_{n}\rangle}+t\frac{s_{n}s_{n}^{T}}{\langle s_{n},y_{n}\rangle}$ is a nonsingular matrix for $t\in(0,\infty)$ and $\langle s_{n},y_{n}\rangle\neq 0$; see [12]. Conducting an eigenvalue analysis on a symmetrized version of $P_{n+1}$, defined by $A_{n+1}=\frac{P_{n+1}+P_{n+1}^{T}}{2},$ we can derive a two-parameter formula for $t$ in the Dai-Liao CG parameter (5):
@@
$d_{n}=\begin{cases}-\nabla f(x_{0}),&\text{ if }n=0,\\ -\theta_{n}\nabla f(x_{n})+\beta_{n-1}d_{n-1}&\text{ if }n\geq 1,\end{cases}$
@@
$z_{n}=y_{n}+h_{n}\|\nabla f(x_{n})\|^{r}s_{n},\quad h_{n}=\nu+\max\left\{-\frac{\langle s_{n},y_{n}\rangle}{\|s_{n}\|^{2}},0\right\}\|\nabla f(x_{n})\|^{-r},$ (7)
with $r,\nu\in(0,\infty)$. The modified secant condition (7) plays a crucial role in establishing the global convergence of the MBFGS method for non-convex functions, see [25].
@@
Based on this conjugate parameter, they construct the following spectral parameter
@@
The primary motivation of this study is to introduce a modified spectral conjugate gradient method and to establish its global convergence for general functions under standard assumptions. This new method consistently satisfies the sufficient descent property, regardless of the line search employed. To demonstrate the effectiveness of the proposed approach, we present numerical results that validate its performance and show its superior convergence compared to existing methods. Furthermore, we explore its application in compressed sensing, emphasizing its practical relevance and significant advantages.
@@ 2 Preliminaries
Let $\mathbb{R}^{m}$ be the Euclidean space and let $\{u_{n}\}$ and $\{v_{n}\}$ be sequences in $\mathbb{R}^{m}$. Let $\{a_{n}\}$ and $\{b_{n}\}$ be sequences in $(0,\infty)$ and $u_{0}\in\mathbb{R}^{m}$ and $b_{0},\sigma\in(0,\infty)$ satisfying
@@
Squaring both side and applying Lemma 1 with $c=-\langle v_{n+1},u_{n+1}\rangle>0$ (from (C2)) and $d=b_{n}|\langle v_{n},u_{n}\rangle|>0$ gives
@@
$\frac{\langle v_{n+1},u_{n+1}\rangle^{2}}{\|u_{n+1}\|^{2}}+\frac{\langle v_{n},u_{n}\rangle^{2}}{\|u_{n}\|^{2}}$ $=\frac{1}{\|u_{n+1}\|^{2}}\left[\langle v_{n+1},u_{n+1}\rangle^{2}+\frac{\|u_{n+1}\|^{2}}{\|u_{n}\|^{2}}\langle v_{n},u_{n}\rangle^{2}\right]$
@@ 3 A modified descent Dai-Liao spectral conjugate gradient method and its convergence analysis
In this section, inspired by the concepts from quasi-Newton methods, we introduce an effective modified spectral conjugate gradient (CG) method.
Motivated by the theoretical properties provided by (7) and the Dai-Liao conjugate gradient parameter (5), we propose the following modified Dai-Liao conjugate parameter:
@@
$t_{n}^{p,q}=p\frac{\|z_{n}\|^{2}}{\langle s_{n},z_{n}\rangle}-q\frac{\langle s_{n},z_{n}\rangle}{\|s_{n}\|^{2}}\text{ for all }n\in\mathbb{N}_{0},$ (18)
@@
Pproposition 6.
Let $f:\mathbb{R}^{m}\to\mathbb{R}$ be a differentiable function and let $\{d_{n}\}$ be a sequence of spectral CG search directions generated by (20). Then we have the following:
@@
    $\langle\nabla f(x_{n}),d_{n}\rangle\leq-\kappa_{n}\|\nabla f(x_{n})\|^{2}$ (21)
@@
    $\kappa_{n}=\left\{\begin{array}[]{ll}1&\text{if }n=0,\\ \theta_{n}-\left(\frac{1}{4p}+|q|\right)&\text{if }n\geq 1.\end{array}\right.$
@@
$\langle\nabla f(x_{n+1}),d_{n+1}\rangle$
$=\langle\nabla f(x_{n+1}),-\theta_{n+1}\nabla f(x_{n+1})+\beta_{n}^{MDDL}d_{n}\rangle$
@@
$\quad-p\frac{\|z_{n}\|^{2}}{\langle s_{n},z_{n}\rangle}\frac{\langle s_{n},\nabla f(x_{n+1})\rangle\langle d_{n},\nabla f(x_{n+1})\rangle}{\langle d_{n},z_{n}\rangle}+q\frac{\langle s_{n},z_{n}\rangle}{\|s_{n}\|^{2}}\frac{\langle s_{n},\nabla f(x_{n+1})\rangle\langle d_{n},\nabla f(x_{n+1})\rangle}{\langle d_{n},z_{n}\rangle}.$
@@
$\langle\nabla f(x_{n+1}),d_{n+1}$
$=-\theta_{n+1}\|\nabla f(x_{n+1})\|^{2}+\frac{\langle\nabla f(x_{n+1}),z_{n}\rangle\langle\nabla f(x_{n+1}),d_{n}\rangle}{\langle d_{n},z_{n}\rangle}-p\frac{\|z_{n}\|^{2}\langle d_{n},\nabla f(x_{n+1})\rangle^{2}}{\langle d_{n},z_{n}\rangle}$
$\quad+q\frac{\langle d_{n},z_{n}\rangle}{\|d_{n}\|^{2}}\frac{\langle d_{n},\nabla f(x_{n+1})\rangle^{2}}{\langle d_{n},z_{n}\rangle}$
$=-\theta_{n+1}\|\nabla f(x_{n+1})\|^{2}+\frac{\langle\nabla f(x_{n+1}),z_{n}\rangle\langle\nabla f(x_{n+1}),d_{n}\rangle}{\langle d_{n},z_{n}\rangle}-p\frac{\|z_{n}\|^{2}\langle d_{n},\nabla f(x_{n+1})\rangle^{2}}{\langle d_{n},z_{n}\rangle^{2}}$
$\quad+q\frac{\langle d_{n},\nabla f(x_{n+1})\rangle^{2}}{\|d_{n}\|^{2}}.$ (22)
@@
${\theta_{n+1}}$
$=\frac{1}{\langle z_{n},\nabla f(x_{n+1})\rangle}(\langle s_{n},\nabla f(x_{n+1})\rangle+\beta_{n}^{MDDL}\langle z_{n},d_{n}\rangle)$
$=\frac{1}{\langle z_{n},\nabla f(x_{n+1})\rangle}\left(\langle s_{n},\nabla f(x_{n+1})\rangle+\left(\frac{\langle\nabla f(x_{n+1}),z_{n}\rangle}{\langle d_{n},z_{n}\rangle}-t_{n}^{p,q}\frac{\langle\nabla f(x_{n+1}),s_{n}\rangle}{\langle d_{n},z_{n}\rangle}\right)\langle z_{n},d_{n}\rangle\right)$
$=\frac{1}{\langle z_{n},\nabla f(x_{n+1})\rangle}\left(\langle s_{n},\nabla f(x_{n+1})\rangle+\frac{\langle\nabla f(x_{n+1}),z_{n}\rangle\langle z_{n},d_{n}\rangle}{\langle d_{n},z_{n}\rangle}-t_{n}^{p,q}\frac{\langle\nabla f(x_{n+1}),s_{n}\rangle\langle z_{n},d_{n}\rangle}{\langle d_{n},z_{n}\rangle}\right)$
$=\frac{1}{\langle z_{n},\nabla f(x_{n+1})\rangle}\left(\langle s_{n},\nabla f(x_{n+1})\rangle+\langle\nabla f(x_{n+1}),z_{n}\rangle-t_{n}^{p,q}\langle\nabla f(x_{n+1}),s_{n}\rangle\right)$
$=1-\frac{1}{\langle z_{n},\nabla f(x_{n+1})\rangle}\left({t_{n}^{p,q}\langle s_{n},\nabla f(x_{n+1})\rangle}-\langle s_{n},\nabla f(x_{n+1})\rangle\right)$ (25)
@@
We denote the $\theta_{n+1}$ obtained in equation (26) by $\theta_{n+1}^{N+}$ and the $\theta_{n+1}$ obtained in equation (27) by $\theta_{n+1}^{N-}$. To obtain the sufficient descent condition and the bounded property of the spectral parameters, we propose the following spectral parameters:
@@
Input: Let $x_{0}\in\mathbb{R}^{n}$ and let $\epsilon,\delta,\sigma,\eta,\tau,r,\nu\in(0,\infty)$ such that $10<\delta<\sigma<$ 1.
@@
Step 2: Determine the step length $\alpha_{n}$ satisfying the strong Wolfe conditions (8) and (9).
@@
where CG parameter $\beta_{n}^{MDDL}$ is given by (19) and spectral parameter is given by (28) or (29).
@@
which implies, from (7) that $h_{n}=\nu$ and hence
$z_{n}=y_{n}+\nu\|\nabla f(x_{n})\|^{r}s_{n},$ (33)
where $r,\nu\in(0,\infty)$.
@@
Let $f:\mathbb{R}^{n}\to\mathbb{R}$ be a function and let $\{x_{n}\}$ be sequence generated by Algorithm 1 such that Assumptions (A1) and (A2) hold. Suppose that $\theta_{n}>\frac{1}{4p}+|q|$ for all $n\in\mathbb{N}$. If $\liminf_{n\to\infty}\|\nabla f(x_{n})\|\neq 0$, then $\sum_{n=0}^{\infty}\frac{\|\nabla f(x_{n})\|^{4}}{\|d_{n}\|^{2}}<\infty.$
@@
Let $u_{n}=d_{n}$, $v_{n}=\nabla f(x_{n})$, $a_{n}=\theta_{n}$ and $b_{n}=\beta_{n}$ in Lemma 3. Since $\theta_{n}>\frac{1}{4p}+|q|$, it follows from Remark 7 that $\langle\nabla f(x_{n}),d_{n}\rangle<0$ for all $n\in\mathbb{N}_{0}$. Also from strong Wolfe condition (9), we have $|\langle\nabla f(x_{n+1}),d_{n}\rangle|\leq-\sigma\langle\nabla f(x_{n}),d_{n}\rangle$. Then, clearly all the conditions (C1), (C2) and (C3) of Lemma 3 is satisfied. Thus, we have
@@
From (36) and Lemma 4, we get $\sum_{n=1}^{\infty}\frac{\|\nabla f(x_{n+1})\|^{4}}{\|d_{n+1}\|^{2}}<\infty.$
@@
Let $f:\mathbb{R}^{n}\to\mathbb{R}$ be a function and let $\{x_{n}\}$ be sequence generated by Algorithm 1 such that Assumptions (A1) and (A2) hold. Then $\liminf_{n\to\infty}\|\nabla f(x_{n})\|=0.$
@@
$\langle d_{n},z_{n}\rangle=\langle d_{n},y_{n}\rangle+\nu\|\nabla f(x_{n})\|^{r}\alpha_{n}\|d_{n}\|^{2}\geq\nu\|\nabla f(x_{n})\|^{r}\alpha_{n}\|d_{n}\|^{2}\geq\nu\xi^{r}\alpha_{n}\|d_{n}\|^{2}.$ (38)
@@
$t_{n}^{p,q}\leq\left|p\frac{\|z_{n}\|^{2}}{\langle s_{n},z_{n}\rangle}-q\frac{\langle s_{n},z_{n}\rangle}{\|s_{n}\|^{2}}\right|\leq p\frac{(L+\nu\gamma^{r})^{2}\|s_{n}\|^{2}}{\mu\|s_{n}\|^{2}}+|q|\frac{(L+\nu\gamma^{r})\|s_{n}\|^{2}}{\|s_{n}\|^{2}}=p\frac{L^{2}}{\mu}+L|q|.$
@@
$\|\beta_{n}^{MDDL}\|$ $=\left\|\frac{\langle\nabla f(x_{n+1}),z_{n}\rangle}{\langle d_{n},z_{n}\rangle}-t_{n}^{p,q}\frac{\langle\nabla f(x_{n+1}),s_{n}\rangle}{\langle d_{n},z_{n}\rangle}\right\|$
@@
$=\left\|\frac{\langle\nabla f(x_{n+1}),z_{n}\rangle}{\langle d_{n},z_{n}\rangle}\right\|+p\left\|\frac{\|z_{n}\|^{2}}{\langle s_{n},z_{n}\rangle}\frac{\langle\nabla f(x_{n+1}),s_{n}\rangle}{\langle d_{n},z_{n}\rangle}\right\|+|q|\left\|\frac{\langle s_{n},z_{n}\rangle}{\|s_{n}\|^{2}}\frac{\langle\nabla f(x_{n+1}),s_{n}\rangle}{\langle d_{n},z_{n}\rangle}\right\|$
@@
where $E=\tau\gamma+\frac{\gamma(L+\nu\gamma^{r})}{\nu\xi^{r}}+\frac{p\gamma(L+\nu\gamma^{r})^{2}}{\nu^{2}\xi^{2r}}+|q|\gamma$. This implies that

### v2

@@ Abstract
Spectral Conjugate Gradient Method , Dai-Liao Method , Optimization , Signal Processing
MSC:
[2020] 65K05 , 90C30 , 49M37
[…]
organization=Department of Mathematics, Banaras Hindu University,city=Varanasi, country=India
[…]
organization=Department of Mathematics, The Technion- Israel Institute of Technology,city=Haifa, country=Israel
[…]
organization=Department of Applied Mathematics and Scientific Computing, Indian Institute of Technology Roorkee,country=India
@@ 1 Introduction
where the stepsize $\alpha_{n}$ is a positive real parameter determined via exact or inexact line search, $x_{n}$ is the previous iterative point, $x_{n+1}$ is the current point and $d_{n}$ is a suitable search direction. The class of algorithms that follow the form (2) is known as line search algorithms. For such algorithms, only the search direction $d_{n}\in\mathbb{R}^{m}$ and the stepsize $\alpha_{n}\in(0,\infty)$ are required. An appropriate descent direction $d_{n}$ must be chosen such that it satisfies the descent condition:
@@
Inspired by Shanno’s matrix perspective on CG algorithms [10], Babaie-Kafaki and Ghanbari [11] observed that the Dai-Liao search directions can be expressed as $d_{n+1}=-P_{n+1}\nabla f(x_{n+1})$ for all $n\in\mathbb{N}_{0}$, where $P_{n+1}=I-\frac{s_{n}y_{n}^{T}}{\langle s_{n},y_{n}\rangle}+t\frac{s_{n}s_{n}^{T}}{\langle s_{n},y_{n}\rangle}$ is a nonsingular matrix for $t\in(0,\infty)$ and $\langle s_{n},y_{n}\rangle\neq 0$; see [12]. Conducting an eigenvalue analysis on a symmetrized version of $P_{n+1}$, defined by $A_{n+1}=\frac{P_{n+1}+P_{n+1}^{T}}{2}$, we can derive a two-parameter formula for $t$ in the Dai-Liao CG parameter (5), given by
@@
$d_{n}=\begin{cases}-\nabla f(x_{0}),&\text{ if }n=0,\\ -\theta_{n}\nabla f(x_{n})+\beta_{n-1}d_{n-1},&\text{ if }n\geq 1,\end{cases}$
@@
$z_{n}=y_{n}+h_{n}\|\nabla f(x_{n})\|^{r}s_{n},\quad h_{n}=\nu+\max\left\{-\frac{\langle s_{n},y_{n}\rangle}{\|s_{n}\|^{2}},0\right\}\|\nabla f(x_{n})\|^{-r}$ (7)
with $r,\nu\in(0,\infty)$. The modified secant condition (7) plays a crucial role in establishing the global convergence of the MBFGS method for non-convex functions; see [25].
@@
Based on this conjugate parameter, they construct the following spectral parameter:
@@
The primary motivation of this study is to introduce a modified spectral conjugate gradient method and to establish its global convergence for general functions under standard assumptions. This new method consistently satisfies the sufficient descent property, regardless of the line search used. To demonstrate the effectiveness of the proposed approach, we present numerical results that validate its performance and show its superior convergence compared to the modified spectral conjugate gradient method of Faramarzi and Amini [26] and the scaled conjugate gradient method of Mard and Fakhari [28]. Furthermore, we explore its application in compressed sensing, emphasizing its practical relevance and significant advantages.
@@ 2 Preliminaries
Let $\mathbb{R}^{m}$ be the Euclidean space and let $\{u_{n}\}$ and $\{v_{n}\}$ be sequences in $\mathbb{R}^{m}$. Let $b_{0},\sigma\in(0,\infty)$ and $u_{0}\in\mathbb{R}^{m}$ and let $\{a_{n}\}$ and $\{b_{n}\}$ be sequences in $(0,\infty)$ satisfying the following conditions:
@@
Squaring both sides and applying Lemma 1 with $c=-\langle v_{n+1},u_{n+1}\rangle>0$ (from (C2)) and $d=b_{n}|\langle v_{n},u_{n}\rangle|>0$ gives
@@
$\frac{\langle v_{n+1},u_{n+1}\rangle^{2}}{\|u_{n+1}\|^{2}}+\frac{\langle v_{n},u_{n}\rangle^{2}}{\|u_{n}\|^{2}}$
$=\frac{1}{\|u_{n+1}\|^{2}}\left[\langle v_{n+1},u_{n+1}\rangle^{2}+\frac{\|u_{n+1}\|^{2}}{\|u_{n}\|^{2}}\langle v_{n},u_{n}\rangle^{2}\right]$
@@ 3 A modified descent Dai-Liao spectral conjugate gradient method and its convergence analysis
In this section, we introduce a modified spectral conjugate gradient (CG) method, inspired by ideas from quasi-Newton methods.
Using the theoretical properties provided by (7) and the Dai-Liao conjugate gradient parameter (5), we propose the following modified Dai-Liao conjugate parameter:
@@
$t_{n}^{p,q}=p\frac{\|z_{n}\|^{2}}{\langle s_{n},z_{n}\rangle}-q\frac{\langle s_{n},z_{n}\rangle}{\|s_{n}\|^{2}}\text{ for all }n\in\mathbb{N}_{0}$ (18)
@@
Proposition 5.
Let $f:\mathbb{R}^{m}\to\mathbb{R}$ be a differentiable function and let $\{d_{n}\}$ be a sequence of spectral CG search directions generated by (20). Then, we have the following:
@@
    $\langle\nabla f(x_{n}),d_{n}\rangle\leq-\kappa_{n}\|\nabla f(x_{n})\|^{2},$ (21)
@@
    $\kappa_{n}=\left\{\begin{array}[]{ll}1,&\text{if }n=0,\\ \theta_{n}-\left(\frac{1}{4p}+|q|\right),&\text{if }n\geq 1.\end{array}\right.$
@@
$\langle\nabla f(x_{n+1}),d_{n+1}\rangle$ $=\langle\nabla f(x_{n+1}),-\theta_{n+1}\nabla f(x_{n+1})+\beta_{n}^{MDDL}d_{n}\rangle$
@@
$\quad-p\frac{\|z_{n}\|^{2}}{\langle s_{n},z_{n}\rangle}\frac{\langle s_{n},\nabla f(x_{n+1})\rangle\langle d_{n},\nabla f(x_{n+1})\rangle}{\langle d_{n},z_{n}\rangle}$
$\quad+q\frac{\langle s_{n},z_{n}\rangle}{\|s_{n}\|^{2}}\frac{\langle s_{n},\nabla f(x_{n+1})\rangle\langle d_{n},\nabla f(x_{n+1})\rangle}{\langle d_{n},z_{n}\rangle}.$
@@
$\langle\nabla f(x_{n+1}),d_{n+1}\rangle$ $=-\theta_{n+1}\|\nabla f(x_{n+1})\|^{2}+\frac{\langle\nabla f(x_{n+1}),z_{n}\rangle\langle\nabla f(x_{n+1}),d_{n}\rangle}{\langle d_{n},z_{n}\rangle}$
$\quad-p\frac{\|z_{n}\|^{2}\langle d_{n},\nabla f(x_{n+1})\rangle^{2}}{\langle d_{n},z_{n}\rangle}+q\frac{\langle d_{n},z_{n}\rangle}{\|d_{n}\|^{2}}\frac{\langle d_{n},\nabla f(x_{n+1})\rangle^{2}}{\langle d_{n},z_{n}\rangle}$
$\quad-p\frac{\|z_{n}\|^{2}\langle d_{n},\nabla f(x_{n+1})\rangle^{2}}{\langle d_{n},z_{n}\rangle^{2}}+q\frac{\langle d_{n},\nabla f(x_{n+1})\rangle^{2}}{\|d_{n}\|^{2}}.$ (22)
@@
${\theta_{n+1}}$ $=\frac{1}{\langle z_{n},\nabla f(x_{n+1})\rangle}(\langle s_{n},\nabla f(x_{n+1})\rangle+\beta_{n}^{MDDL}\langle z_{n},d_{n}\rangle)$
$=\frac{1}{\langle z_{n},\nabla f(x_{n+1})\rangle}\left(\langle s_{n},\nabla f(x_{n+1})\rangle+\left(\frac{\langle\nabla f(x_{n+1}),z_{n}\rangle}{\langle d_{n},z_{n}\rangle}\right.\right.$
$\quad\left.\left.-t_{n}^{p,q}\frac{\langle\nabla f(x_{n+1}),s_{n}\rangle}{\langle d_{n},z_{n}\rangle}\right)\langle z_{n},d_{n}\rangle\right)$
$=\frac{1}{\langle z_{n},\nabla f(x_{n+1})\rangle}\left(\langle s_{n},\nabla f(x_{n+1})\rangle+\frac{\langle\nabla f(x_{n+1}),z_{n}\rangle\langle z_{n},d_{n}\rangle}{\langle d_{n},z_{n}\rangle}\right.$
$\quad\left.-t_{n}^{p,q}\frac{\langle\nabla f(x_{n+1}),s_{n}\rangle\langle z_{n},d_{n}\rangle}{\langle d_{n},z_{n}\rangle}\right)$
$=\frac{\langle s_{n},\nabla f(x_{n+1})\rangle+\langle\nabla f(x_{n+1}),z_{n}\rangle-t_{n}^{p,q}\langle\nabla f(x_{n+1}),s_{n}\rangle}{\langle z_{n},\nabla f(x_{n+1})\rangle}$
$=1-\frac{{t_{n}^{p,q}\langle s_{n},\nabla f(x_{n+1})\rangle}-\langle s_{n},\nabla f(x_{n+1})\rangle}{\langle z_{n},\nabla f(x_{n+1})\rangle}$ (25)
@@
We denote the $\theta_{n+1}$ obtained in equation (26) by $\theta_{n+1}^{R}$ and the $\theta_{n+1}$ obtained in equation (27) by $\theta_{n+1}^{N}$. To obtain the sufficient descent condition and the bounded property of the spectral parameters, we propose the following spectral parameters:
@@
Input: Let $x_{0}\in\mathbb{R}^{n}$ and let $\epsilon,\delta,\sigma,\eta,\tau,r,\nu\in(0,\infty)$ such that
     $10<\delta<\sigma<$ 1.
@@
Step 2: Determine the step length $\alpha_{n}$ satisfying the strong Wolfe
    conditions (8) and (9).
@@
      where CG parameter $\beta_{n}^{MDDL}$ is given by (19) and spectral
    parameter is given by (28) or (29).
@@
which, from (7), implies that $h_{n}=\nu$ and hence
$z_{n}=y_{n}+\nu\|\nabla f(x_{n})\|^{r}s_{n}.$ (33)
@@
Let $f:\mathbb{R}^{n}\to\mathbb{R}$ be a function and let $\{x_{n}\}$ be the sequence generated by Algorithm 1 such that Assumptions (A1) and (A2) hold. Suppose that $\theta_{n}>\frac{1}{4p}+|q|$ for all $n\in\mathbb{N}$. If $\liminf_{n\to\infty}\|\nabla f(x_{n})\|\neq 0$, then $\sum_{n=0}^{\infty}\frac{\|\nabla f(x_{n})\|^{4}}{\|d_{n}\|^{2}}<\infty.$
@@
Let $u_{n}=d_{n}$, $v_{n}=\nabla f(x_{n})$, $a_{n}=\theta_{n}$ and $b_{n}=\beta_{n}$ in Lemma 3. Since $\theta_{n}>\frac{1}{4p}+|q|$, it follows from Remark 6 that $\langle\nabla f(x_{n}),d_{n}\rangle<0$ for all $n\in\mathbb{N}_{0}$. Also from strong Wolfe condition (9), we have $|\langle\nabla f(x_{n+1}),d_{n}\rangle|\leq-\sigma\langle\nabla f(x_{n}),d_{n}\rangle$. Then, clearly all the conditions (C1), (C2) and (C3) of Lemma 3 are satisfied. Thus, we have
@@
From (36) and Lemma 4, we get that $\sum_{n=1}^{\infty}\frac{\|\nabla f(x_{n+1})\|^{4}}{\|d_{n+1}\|^{2}}<\infty.$
@@
Let $f:\mathbb{R}^{n}\to\mathbb{R}$ be a function and let $\{x_{n}\}$ be the sequence generated by Algorithm 1 such that Assumptions (A1) and (A2) hold. Then $\liminf_{n\to\infty}\|\nabla f(x_{n})\|=0.$
@@
$\langle d_{n},z_{n}\rangle$ $=\langle d_{n},y_{n}\rangle+\nu\|\nabla f(x_{n})\|^{r}\alpha_{n}\|d_{n}\|^{2}$
$\geq\nu\|\nabla f(x_{n})\|^{r}\alpha_{n}\|d_{n}\|^{2}\geq\nu\xi^{r}\alpha_{n}\|d_{n}\|^{2}.$ (38)
@@
$t_{n}^{p,q}$ $\leq\left|p\frac{\|z_{n}\|^{2}}{\langle s_{n},z_{n}\rangle}-q\frac{\langle s_{n},z_{n}\rangle}{\|s_{n}\|^{2}}\right|$
$\leq p\frac{(L+\nu\gamma^{r})^{2}\|s_{n}\|^{2}}{\mu\|s_{n}\|^{2}}+|q|\frac{(L+\nu\gamma^{r})\|s_{n}\|^{2}}{\|s_{n}\|^{2}}=p\frac{L^{2}}{\mu}+L|q|.$
@@
$\|\beta_{n}^{MDDL}\|$
$=\left\|\frac{\langle\nabla f(x_{n+1}),z_{n}\rangle}{\langle d_{n},z_{n}\rangle}-t_{n}^{p,q}\frac{\langle\nabla f(x_{n+1}),s_{n}\rangle}{\langle d_{n},z_{n}\rangle}\right\|$
@@
$=\left\|\frac{\langle\nabla f(x_{n+1}),z_{n}\rangle}{\langle d_{n},z_{n}\rangle}\right\|+p\left\|\frac{\|z_{n}\|^{2}}{\langle s_{n},z_{n}\rangle}\frac{\langle\nabla f(x_{n+1}),s_{n}\rangle}{\langle d_{n},z_{n}\rangle}\right\|$
$\quad+|q|\left\|\frac{\langle s_{n},z_{n}\rangle}{\|s_{n}\|^{2}}\frac{\langle\nabla f(x_{n+1}),s_{n}\rangle}{\langle d_{n},z_{n}\rangle}\right\|$
@@
where $E=\tau\gamma+\frac{\gamma(L+\nu\gamma^{r})}{\nu\xi^{r}}+\frac{p\gamma(L+\nu\gamma^{r})^{2}}{\nu^{2}\xi^{2r}}+|q|\gamma$. Thus,